Runplan of the Day

From NPDGamma Wiki

Jump to: navigation, search

The shift worker is NOT to take it upon her/himself to make decisions. If you think we need to change the mode of running, call a senior experiment contact (Seppo or Nadia). IF there's a problem --> Do NOT wait for it to magically fix itself - call the appropriate person on the list.


Now that there's no one on the owl shift, make note to do the following:

  • When leaving the swing shift, make sure clover and the relevant directory on basestar each have at least 24GB of free space for the owl shift data
  • When arriving on the day shift, go through the owl data and make sure it's all good.

Contents

August 13, 2013

The beam is off since 8:00 am. It will be back around 11:00 am according to IHCs. The polarimetry set-up is done. Once the beam is back, we will be doing the polarimetry. The plan after polarimetry is to place the Cl target inside the SF and take Cl data.

For the evening shift, once the chlorine target has been removed from the cave we will begin regular hydrogen production runs. At this point we are very close to being fully converted. Take one or two pedestal runs before opening the shutter, and then after the shutter is open it can be left open for the remaining couple hours of the shift with the DAQ running in series mode. It will be left in this configuration taking series data overnight. We will continue to watch for any further conversion using the M4 data.

Plan for data-taking while target is converting (Aug 12,2013)

We are taking data runs about every hour and a half to look at the ortho-para conversion using M4. Once every hour and a half or so, take one pedestal run, two or three data runs, and then a pedestal run. Leave the shutter closed when not taking data runs to not unnecessarily irradiate the apparatus. Please check to make sure that every data run is good by looking at the pedestal-subtraction M4/M1 ratio. We want to plot this quantity versus time to watch the conversion. We have developed an analysis routine on clover to make this more straightforward. At the end of the swing shift the DAQ should be left running taking data overnight with shutter open.

Plan for beam down time from 3rd to 6th Feb.

Since we do not have beam but target is running we have decided that we do not need to baby sit the target. The experts will follow target remotely. seppo

A plan for Jan 15 and Jan 16

We will take transmission runs periodically throughout the swing shift of Jan 15th. At midnight before leaving the swing shift we will start a series of shutter closed noise runs. Shutter closed noise runs will continue into the morning shift of the 16th. Do not disturb the shutter in the morning until notified from someone such as Nadia or Seppo.

A plan for Wed Jan-09

A plan could be to continue tgt filling and then afternoon when Zhaowen is in take - if not needed any more - the Cl tgt out and insert the CsI tgt that we can then run over nite again.


Draft action plan till 4pm or when beam goes away

Yesterday, Wed, we tuned the OPC monitor to max signal. Today we should use beam time for noise study/reduction and backgrounds.

  • Noise study/reduction; since we stop hydrogen runs, we can start to turn off and disconnect components that we do not need in cross section measurement, such as field and SF which have couplings to OPC monitor.
  • Background study with 60Hz; using blanks, absorbers, and collimators we can learn the neutronic and electronic backgrounds with 60 Hz.
  • The 1st part of the drift measurement. For this we need to use low temp coefficient resistors 50-100 ppm/C and precision pulser. We could have also this working today and could be done after we loose 60Hz beam. The system needs to be put together and tested in bench first

Draft Action Plan for December Total Cross Section Measurement

Link to proposed run plan that was discussed in an analysis meeting a few weeks ago. [1]

Assumption is 60 Hz beam on till Thu Dec-20 at 4pm, then there is a 8-h shift of 10Hz beam. ( NF: Not exactly. This is still under discussion and negotiation)

Action plan for getting ready:

  • 1. On Tue 18th techs will mount support structure on beam stop. This will take about 4h. We try to get started with them in morning, this is possible if the fittings have been arrived.
  • 2. Zhaowen and seppo will assemble system on table as much as possible and then install the system in beam. This could take place on Tue afternoon.
  • 3. After installation they will continue with reduction of noise.
  • 4. Then we can start tuning with beam; with a few scans to max signal then check pedestal have they been reduced.
  • 5. Wednesday is the contingency day.

Thoughts for run Assumptions are that

  • M4 already tuned with beam
  • pedestal is 10% from signal as it is now - it should be smaller
  • choppers parked open
  • we aim better than 3% measurement cross section

Run plan:

  • 1. start with pedestal need to be measured better than 10%
  • 2. a few background runs to learn what backgrounds are
  • 3. cross section runs as many as possible
  • 4. pedestal
  • 5. background runs - if backgrounds are small we do not need too many runs
  • 6. pedestal

Problems:

  • how to measure longer wavelength component under signal, with 10 Hz we talk about 20 angstrom or longer wavelength neutrons
  • how to do a proper background measurement; M4 will be at the same location as earlier. We block the beam on the cryostat, we can block the beam on the front of the tube, we can block the beam on the detector entry.
  • how to establish the drift of our system. When we coming back on Jan with the empty target the system has drift and we need to have a way to measure it and do a possible correction.

December 3rd-onwards, 2012

Take data. No more decays for a while.


December 1-2nd, 2012

Take data.

Do a shutter closing decay measurement during swing shift with stable beam - i.e. after beam has been on for at least a few minutes and been stable, close the shutter in the beginning of a run. Continue taking data for one hour. Reopen the shutter. Make notes.

November 29-30, 2012

Take data.

Do a shutter closing decay measurement during swing shift with stable beam - i.e. after beam has been on for at least a few minutes and been stable, close the shutter in the beginning of a run. Continue taking data for one hour. Reopen the shutter. Make notes.

November 28, 2012

Take data whenever beam finally comes back.

Do a shutter closing decay measurement early on, but with stable beam - i.e. after beam has been on for at least a few minutes and been stable, close the shutter in the beginning of a run. Continue taking data for one hour. Reopen the shutter. Make notes.

November 27, 2012

Remove Chlorine target after verifying data.

Load hydrogen gains, take hydrogen data.

Do a shutter closing decay measurement during the swing shift - i.e. after beam has been on for at least a few minutes and been stable, close the shutter in the beginning of a run. Continue taking data for one hour. Reopen the shutter. Make notes.

September 25, 2012

Stop taking LED runs with #94586. Triangle 2Vpp 20Hz connected to M2 input in VME3 to see, maybe, what is going on when we have bad spin sequence.

September 22, 2012

After the failure of the mercury target last night Elise started to take noise runs. A plan is to continue with the noise runs till Sunday morning. Michael will then switch the LEDs on and then we need to run about two days with LEDs. After that we need to think if there is some tests that we want to do, for instance, we could run for a while with our our T0 and see if we have the same spin sequence failing rate as we have seen. This would test if the T0 provided by the facility has still problems. Also this would test if the daq timing from the facility introduces noise or reduces noise.


September 16th-17th-18th, 2012

According to IHC, the beam will be off starting at 8:00 am on September 16,2012. Matthew is on owl shift on September 16,2012. So, Matt: Call IHC at 4:00 am to confirm that the beam will be off at 8:00 am. Once you confirm the time being 8:00 am, close the shutter during a run at 5:30 am to take decay data. We need shutter closing and decay data. This is extremely important decay data, because we do not have the background calculations for Ring 1. Make sure that on the comment line it says " decay data". Change the comment to " Noise data" at 8:00 am before you leave. We will be taking noise data for noise asymmetry measurements. Once Kyle checks the data and talks to Seppo, Seppo will come and turn the LEDs on for the LED measurements.

September 6th, 2012

Background measurements

Chlorine

Monitor new T0

September 5th, 2012

Take noise data to see if the timing computer reboot has fixed the occasional missing T0 problem.

September 4th, 2012

Continue taking transmission data

June 16-17, 2012

They decided to run to the end of cycle with 400-kW beam, which is useless for us for H2 runs or any serious studies. At present - if somebody comes up with a great idea what to do with the beam - we do not run shutter open. The tgt folks are thinking and will start tgt testing tomorrow. Parallel we will do false asymmetry runs without beam injecting charge to M2 and first running without LEDs.

June 15-16, 2012

Keep running with the LH2 target till we are ready with the Bessel circuit. The Bessell filter/amplifier board for one ring has left LANL today, it should arrive to lab tomorrow and we should have it afternoon. We also have the NIM box tooled, it requires still powering that we will get working by afternoon. This means that we can start bench testing the circuit first and then do testing with detectors and DAQ and then if everything works we can start running with beam.


June 13-14-15, 2012

LANL is struggling with the Bessel filter/amplifier board, test showed that there is a short in the circuit. It can be that the board will arrive at the last moment=during the week end, we will loose beam on Monday morning. We need to take care that we will receive the board even if it is the weekend. Serpil has ordered the empty NIM boxes and I "borrowed" one old used NIM box from Orela's collection. The back and front panels of this box are in work shop to be cut for the board. Now we are working with powering the circuit. At the end of the week everything should be ready for testing of the one ring with the new circuit. Till then we keep collecting the LH2 data.

June 11-12, 2012

Since there is no break in production beam delivery, run till 1pm with Al, then close shutter and let Al to cool about 2 hours during cool period take data - day runs. At 3 pm Zhaowen and seppo will take the Al tgt out. Then run with LH2 till we get the Bessel filter/amplifier circuit and NIM box ready which should take place on Wednesday.

June 09-10, 2012

Continue taking Aluminum data.

June 08, 2012

After verifying the accumulated statistics on Chlorine on the morning of June 07th, the plan is to test the aluminum target. This plan is contingent on having the aluminum target ready. If it is, we will install it at the back of the spin flipper, and confirm the beam polarization with the polarized 3he cell. If that checks out and the SF current is okay, we install additional shielding on the backside of the aluminum target and take data until 1e-7 is achieved or Monday comes. If we end up depolarizing the beam, then we revert to running hydrogen.

RUNNING WITH ALUMINUM - SF is okay, Polarization is okay, signal size is small. Keep running until we have enough statistics to see if this is worthwhile.

June 07, 2012

Since there was no beam after 9:30 a.m. on June 06, we are still one day short of needed chlorine statistics. So, when beam returns, continue running with chlorine.

June 05-06, 2012

Load the Cl tgt and start to take data as soon as beam is back. Keep running Cl as long as Nadia has enough statistics or and we have Al ready to be used.

June 03-04, 2012

Keep DAQ running; after beam was turned off do 2 h decay part - done - then keep running, as additive noise runs without beam till tomorrow morning. From data we can study the level of electronic asymmetry and the stability of the system, we have seen slow changes in the long decays. Is this level of stability of the system; this doesn't effect on the gamma asymmetry, we cancel drifts to 2nd order but, slowly varying signal gives performance limit of our system. Tomorrow morning we can switch to LED electronic asymmetry runs and then if the Bessel filter amplifiers arrive on Monday we can test them withoout beam. On Monday we will take the Al tgt holder/support to machine shop.


June 02-03, 2012

Continue running with hydrogen till the maintenance break. According to SNS schedule the 2-3 day maintenance break will start in Sunday morning June 3rd and then beam should be back in production in morning June 6th. At that time we should have in hand the Cl and Al targets and also the Bessel filter. Since new development with interpreting the Iodine decay, it can be that we do not need the CsI target, the order is on hold. Before the end of the maintenance break we need to decide how to go forward.

June 01, 2012

Continue running with hydrogen. Tomorrow, Friday, we should have a better idea what can we do, we are limited with targets we have Cl tgt in hand but we need to cut significantly beam to keep dose levelsoutside cave acceptable. We will get tomorrow teflon to build the Al tgt, we will have target ready on Monday. CsI material we have not yet ordered.

May 31, 2012

Continue running with hydrogen until a plan for ending of the run cycle has been fully developed. If during the run there is beam off moments, lasting hour or so, keep taking data. These are good decay measurements in a case that the beam has been stable before the beam off. Kyle is calculating dose rates from our targets, Al, Cl, CsI before we proceed with design and fabrication. The CCl4 target we have in hand, issue here is how much beam can interact with it without exceeding allowed rad levels, how much collimation needed. The thin Al target needs a new holder so that we can effectively wrap it with Li shielding, Al has significantly larger scattering cross section than absorption. For the CsI tgt material and holder material need to be ordered. We can start with the Cl tgt as soon as we know the beam size from Kyle.


May 30, 2012

Continue running with hydrogen until a plan for ending of the run cycle has been developed.

May 29, 2012

run with hydrogen until maintenance break. Continue running with hydrogen after the maintenance break.

May 26, 2012

Switch LEDs off at around 10 a.m. (Zhaowen will come in for this). Take data for 12 hours with LEDs OFF beam on

Keep the beam on (still blocked, M1 on). Gains will not require changing.

At this point, close the secondary shutter at the beginning of a run. Continue to take data for 1 hour.

Resume hydrogen data taking (load May25 gains). M1 on or off based on the outcome of LED measurement.

May 25, 2012

By 10 a.m., switched to beam blocked configuration. M1 is running on M4's battery, plugged into VME3, signal looks good. LEDs on (20 is saturated), beam blocked.

WILL RUN WITH LEDS until 10 a.m. on May 26, 2012

May 24, 2012

midnight - 11a.m. -- Data taking with beam blocked, LEDs on, M1, M2 unplugged

11 a.m. -- M4 (triple monitor) also unplugged, data taking with beam blocked, LEDs on, NO monitors at all, continues.

-->Will look at asymmetry after 80 runs. Decision will be made at that point about what to do next.

Personal tools